HIGH RESOLUTION INFRARED SPECTROSCOPY OF WATER MONOMERS AND WATER CLUSTERS IN SOLID PARAHYDROGEN

MARIO E. FAJARDO, AFRL/MNAV, Flight Vehicles Integration Branch, Assessment and Demonstrations Division, U.S. Air Force Research Lab, 101 West Eglin Blvd., Suite 326, Eglin AFB, FL 32542-6810. mario.fajardo@eglin.af.mil.

I report on progress made in assigning the various infrared (IR) absorption features observed in the spectra of water-doped cryogenic molecular hydrogen solids. Since water is a ubiquitous impurity in our samples, this ”cataloging” stage is a prerequisite to correctly identifying the IR absorptions due to other dopant species of interest. The water assignments are complicated by a number of phenomena, including: (1) clustering of H_{2}O molecules, (2) clustering of residual ortho-H_{2} molecules with water molecules and clusters, and (3) the surprisingly rapid ortho/para conversion of H_{2}O and D_{2}O molecules, both isolated and in clusters. A few highlights: (a) H_{2}O, D_{2}O, and HDO monomers all exist as very slightly hindered rotors in para-H_{2} solids, (b) clustering of these monomers produces novel non-minimum-energy water cluster configurations such as the cyclic water hexamer reported in J. Chem. Phys. v115, p6807 (2001), (c) clustering of water monomers with residual ortho-H_{2} yields spectra of ”non-rotating” water molecules and as-yet un-assigned complex spectra presumably due to ortho-H_{2}/water van der Waals clusters, and (d) the ortho-to-para H_{2}O and para-to-ortho D_{2}O conversion of isolated monomers both follow single-exponential decay profiles with time constants of 1900 s and 860 s, respectively.